20% of all customers subscribe to phone service.

70% of all customers subscribe to internet service.

5% of all customers subscribe to only phone service.

Let A be the event that a customer subscribes to phone service.

Let B be the event that a customer subscribes to internet service.

Of those individuals who subscribe to at least one of the two services, what proportion subscribe to both services. (round to three decimal places.)
Answer #1

P(A) = 0.2

P(I) = 0.7

P(A) - P(A and I) = 0.05

P(A and I) = 0.2 - 0.05 = 0.15

P(A or I) = P(A) + P(I) - P(A and I) = 0.2 + 0.7 - 0.15 = 0.75

P(at least one of the services) = 0.750

P(both services) = 0.150